OpenStudy (anonymous):

Derive this equation:

OpenStudy (anonymous):

\[y=9\sqrt{\cos(x/3)}+2\]

OpenStudy (anonymous):

\[D(y)=9(1/2) \left( \sqrt{\cos(x/3)}\right)^{-1}(-1/3)\sin(x/3)= \frac{-3 \sin(x/3)}{2\sqrt{\cos(x/3)}}\]

OpenStudy (anonymous):

Thanks for your reply! Why did you raise the radical term to -1? I raised it to -(1/2) in my original solution.

OpenStudy (anonymous):

It is the same I have (()^1/2)^-1 using (a^b)^c = a^(bc) you get the same thing.

OpenStudy (anonymous):

Ooh ok, so you wrote it differently! I just had some errors when I simplified the expression. Thank you so much!